what is the sum or difference identity for the follow:
sin 165

Answer by jsmallt9(3759)

With the 5 at the end of 165, you should be thinking that a multiple of 45 would be involved (because all the other special angles end in 0). Here are several possibilities:
sin(165) = sin(120+45) = sin(120)*cos(45) + sin(45)*cos(120)
sin(165) = sin(135+30) = sin(135)*cos(30) + sin(30)*cos(135)
sin(165) = sin(210-45) = sin(210)*cos(45) - sin(45)*cos(210)